OpenStudy (anonymous):

Find the 5 numbers that have a mean of 6, a median of 3, and a range of 19.

OpenStudy (amistre64):

1 to 20 = 19 2 to 21 = 19 3 to 23 = 19 try; 1 b 3 d 20 1+3+20 + b +d --------------- = 6 5 b+d = 30 - 24 = 6; 2 and 4, 3 and 3 ?? 1,2,3,4,20; has a mean of 6, has a median of 3, and a range of 19

OpenStudy (anonymous):

Thank you so so much. I hope someone does a really nice thing for you very soon